OpenStudy (anonymous):

Simplify the expression. 1/2(-sin2x - cos 2x)

OpenStudy (anonymous):

\[\Huge {1 \over 2(-\sin2x-\cos2x)}\] Os it this?

OpenStudy (anonymous):

\[\Huge {1 \over 2(-\sin^2x-\cos^2x)}\] Or this?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } (-\sin2x - \cos 2x) \]